Circular No. 6632

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970402 AND BL CIRCINI T. E. Harrison, B. J. McNamara, and P. A. Mason, New Mexico State

University (NMSU), for the BATSE/COMPTEL/NMSU Rapid Response Network (McNamara et al. 1996, Ap.J. Supp. 103, 173), report: “We have examined a U.K. Schmidt Telescope (UKST) plate obtained on Apr. 3.73 UT by F. G. Wilson (Anglo-Australian Observatory) at the position of the reported BeppoSAX x-ray counterpart for GRB 970402 (IAUC 6617). No sources within the error circle appear to have exhibited significant changes in brightness when compared to the ESO Survey plates. We confirm the observations (IAUC 6628) that BL Cir is very faint. On the new UKST plate, with a bandpass centered at 590 nm, BL Cir is at V ∼ 20. On the ESO blue and red plates, BL Cir is within its ‘normal’ brightness range of m_{pg} ∼ 15–16. BL Cir has been classified as a long-period variable (Hoffmeister 1965, A.N. 289, 1). The dramatic fading, along with the G–K spectral type suggest that BL Cir may be a previously unknown R CrB star. Follow-up infrared observations would be useful in confirming this hypothesis.”

SUPERNOVA 1997bs IN NGC 3627 M. Cavagna reports the following precise position for SN 1997bs, ob-

tained at Sormano with the 0.5-m reflector + CCD by F. Manca and himself on Apr. 17.850 UT: α = 11^{h}20^{m}14^{s}.25, δ = +12^{o}58′19′′.6 (equinox 2000.0).

X2127+119 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; A. Peele, GSFC and National Research Council; and D. A. Smith, Massachusetts Institute of Technology (MIT), report on behalf of the RXTE ASM team at GSFC/MIT: “Observations of this low-mass x-ray binary in M15, obtained between 1996 Jan. 1 and 1997 Apr. 10 with the ASM on RXTE, show (in addition to considerable non-periodic variability) the presence of a modulation with a period of 37 days about the average 2–12-keV flux of 15 mCrab. This modulation is most prominent in the lowest energy band (1.3–3.0 keV) at an average level of ∼ 35 percent and a FWHM in the Fourier transform of ∼ 3.5 days, but is undetectable in the highest energy band (4.8–12 keV). Maximum brightness occurs at an epoch of ∼ 1996 May 25. Modulation at the 17.1123-hr orbital period is also seen. We encourage continued observations at x-ray and optical wavebands to search for other manifestations of the 37-day period and to determine whether this is a persistent property of X2127+119.”

1997 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green